OpenStudy (anonymous):

Solve for x: |2x + 6| − 4 = 20 x = 9 and x = 11 x = −9 and x = 15 x = 9 and x = −15 No solution

OpenStudy (anonymous):

im far behind being far behind i need majore help please :(

OpenStudy (campbell_st):

start by adding 4 to both sides of the equation and you get |2x + 6| = 24 since its an absolute value, you will need to solve 2 equations 2x + 6 = 24 and 2x + 6 = -24 just solve both for x

OpenStudy (anonymous):

mkay ill try

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

still don't get this :((

OpenStudy (campbell_st):

you just need to practice... and with absolute value questions like this just get it to the absolute value is equal to a number then solve the positive and negative version

OpenStudy (anonymous):

theres no solution right

OpenStudy (campbell_st):

there are 2 solutions subtract 6 from both sides of each equation 2x = 18 and 2x = -30 now divide each equation by 2 to get the answers

OpenStudy (anonymous):

ohhhh mkay thanks very much man :)